• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name rohe

    Log entry time 00:04:26 on March10,2006

    Entry number 166052

    keyword=swing shift summary

    continue data taking on carbon. Stopped at 16:34 for Moeller measurement. We have collected about 45 min at 2 muA on carbon --> completed.

    20:00 Preparing fro controlled access for repairing the veto of neutron detector.

    22:15: work done in hall

    Beam tune until 23:00.

    starting on H2.

    23:00 target pol. measurements
    2825 short run
    2826: BB threshold total shower set to -120 mV